INFRARED ABSORPTION SPECTRUM OF THE SIMPLEST CRIEGEE INTERMEDIATE CH_{2}OO

YU-TE SU, YU-HSUAN HUANG, HENRYK WITEK, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan; \underline{YUAN-PERN LEE}, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

The Criegee intermediates are carbonyl oxides postulated to play key roles in the reactions of ozone with unsaturated hydrocarbons; these reactions constitute an important mechanism for the removal of unsaturated hydrocarbons and for the production of OH in the atmosphere. Here we report the transient infrared (IR) absorption spectrum of the simplest Criegee intermediate CH_{2}OO, produced from CH_{2}I + O_{2} in a flow reactor, using a step-scan Fourier-transform spectrometer coupled with a mulitpass absorption cell. CH_{2}I was produced on photolysis of CH_{2}I_{2} with laser light at 248 nm. The five bands observed near 1435, 1286, 1241, 908, and 848 cm^{−1} provide definitive identification of this intermediate on comparison of anharmonic vibrational wavenumbers and relative IR intensities predicted using quadratic force field obtained with the NEVPT2 method implemented in the MOLPRO quantum chemistry package using the CASSCF(8,8) reference wave function. Observed rotational contours also agree with those simulated based on rotational constants predicted with the B3LYP method. Observation of the out-of-plane CH_{2}-wagging (ν_{8}) mode near 848 cm^{−1} with a characteristic c-type structure showing a prominent Q-branch further supports the assignment of the observed features to a planar CH_{2}OO. The observed vibrational frequencies are more consistent with a zwitterion rather than a diradical structure of CH_{2}OO. The direct IR detection of CH_{2}OO should prove useful for kinetic and mechanistic investigations of the Criegee mechanism.